how do i find the approximate radius of a sphere with a surface area of 40 pie square feet?

Answer by stanbon(75887)

how do i find the approximate radius of a sphere with a surface area of 40 pie square feet?
----
SA = 4(pi)r^2
----
40(pi) = 4(pi)r^2
10 = r^2
radius = sqrt(10) = 3.1623 ft.
=============================
Cheers,
Stan Hanson
=============================